VIS Meeting on 2015/11/27
2015/11/27 13:30 - 15:40
Kamioka NAOJ Room + SeeVogh
Participants: Takahashi, Shoda, Fujii, Ohishi, Fabian, Aso, Mark, Hirata, Okutomi, Tatsumi
- Test Hanging Status (Takahashi)
- IMRM was set up. OSEMs were installed and cables were connected.
- Signals from IMRM and TM can be obtained in the digital system
- Now we have a problem with the digital system: Signals can be seen with EPICS, but not with NDS and DAQ system. We don't know when this problem will be fixed.
- Picomotors: a driver and cables are there but need to be connected.
Fishing rod: not yet connected. Takahashi will bring the driver from TAMA on Tue.
LVDT: connected to the driver but not working -> Takahashi will bring a spare coil from Mitaka to check the driver circuit outside the tunnel on Tue. Who will do this ? -> Talk with Miyakawa-san (Aso).
A cable clamp to be attached to the suspension rod has a wrong size collet. Fabian will bring the proper collets from Mitaka. -> Will get it on Monday next week.
- Payload weight issue
- BF load check at the ATC: BF#1 was sent back to Mitaka. Takahashi, Hirata, Saito (ATC) will check the load on 12/3.
- Digital system (Okutomi, Fujii)
- All the OSEM signals are now visible in the digital system.
- Adding modules for LVDTs.
- Traverser actuation check (Shoda)
- Traverser cables were all made but not connected.
- Traverser stop mechanism
- New designs of stoppers for traverser came from Uraguchi-san (ATC) to allow us to stop traverser at any point.
Shoda will check the proposed designs and make a recommendation to the VIS members.
- Clean work procedure (Aso)
Particle count record: Shoda collected data, Need calibration. Shoda will ask Miyoki-san.
- PR2 preparation (Mark, Fabian)
- We don't have the temporary clean booth now.
- Everything brought into the PR2 area are now covered with clean cling films.
- The PR2 floor is still dirty. The cleaning company will come on Monday. We can ask them to clean the floor again if we wish.
- Mark is ready to get a driver's licence
- PR CAD model and assembly drawings (Hirata)
Done: Cable clamps on the BF. Short jacks for Type-Bp inner frame. -> 3D CAD model is complete.
- To Do: 2D assembly drawings
Need to give the rotation angle of the traverser top plate (Aso)
- Cable status (Takahashi)
Made a cable list -> Put a link to the excel file on Wiki. Takahashi will improve the table.
- Call for shift for cable checking
Ask the student if he can come on the week of 12/7 (Takahashi) -> He can come on the week of 12/7.
- Additional PR mirror box (Tatsumi):
- Delivered on 11/27
- PR2 gluing schedule (Tatsumi)
- In order to expedite the PR2 installation, we want to glue the flags before the first contact peel-off in the hanging test.
- The MC first contact will be peeled off in the next week
Coordinate with relevant people (Hirose, Tatsumi, Takahashi, Ishizaki) (Aso)
- Production of additional PR winches (Akutsu)
- RM compensation mass
Design status (Fabian): Will finish it after the ELITES meeting by the end of the next week.
Order status (Takahashi): Placed an order. Will check the delivery date.
- RM EQ stop screw position (Hirata)
Out sourcing (Takahashi): Placed an order. Need to send the plates to the company.
- PR Optical Bench (Akutsu)
- OSEM assembly status (Akutsu)
- Rails for the hanging frame (Hirata)
Hirata will finish exchanging the balls and send them to Kamioka by the end of the next week
- Next week (from 11/24) plan
- Test hanging
- Tue: Takahashi will bring a LVDT coil and a stepping motor driver. Check the fishing rod actuation.
- Fri: VIS meeting @ Mitaka
- PR2 preparation work
- Mon: Clean PR2
- Tue: Cleaning and preparation
- Wed: Alignment of the frames
- Thu: Put the IM on the frame
- Fri: VIS meeting @ Mitaka
- Test hanging
- Travel (Week of 11/30)
- Takahashi: Tue
- Shoda: no
- Mark: Mon(aft) - Thu
- Fabian: no
- Okutomi: Mon-Tue
- Fujii: no
- Ishizaki: Mon - Thu
- Hirata: no
- Ohishi: Mon - Fri
- Travel (Week of 12/7)
- Takahashi: Mon - Fri
- Shoda: Mon - Fri
- Mark: Mon - Fri
- Fabian: Mon - Fri
- Okutomi: Mon - Fri
- Fujii: Mon - Fri
- Ohishi: Tue - Fri
- CLIO BS
- EO approved the plan
- No VIS man power required for this work
- What to do with BS after iKAGRA
- After iKAGRA the BS chamber will have three vacuum tubes connected.
- Is it a viable option to install SR2 first ?
- We need some kind of hanging test for Type-B system.
- Probably doing the test with BS is the best because we can test both the payload part (different from PR) and the upper part (SF and TF). Then we probably do not need a separate hanging test for SR because the payload part of SR is the same as PR.
- We need to decide on the plan by the end of Dec.
- BS Assembly Frame Design (Aso, Hirata)
- BS flag design (Mark, Fabian)
- BS gluing jigs (Tatsumi)
- Not enough crane height problem (Hirata, Saito)
- IP base clamp (Fabian)
- Bread board spring (Takahashi)
- BS clean bag (Aso)
- ETM installation procedure (Aso)
- Aso sent a compiled information to Mirapro on 11/26, but it is not yet complete.
Takahashi will send Mirapro information about the rails
Hirata will send Aso the ball exchange procedure
- Hirata did a stress analysis of the jack arms: Al is probably fine but for safety we decided to change the material to SUS
- Watch dog system (Tanaka)
- Review next week
- GAS blade inspection by Virgo (Takahashi)
Waiting for shipping information: Takahashi-san will send Ando-san a reminder -> not yet.
- Request to the ATC for the next year
- We need to tell ATC our requests for the next fiscal year.
- RM production already started.
Production of the blades of the Type-A BF: We should check the consistency of the optimal load at the TAMA prototype experiment and G&M measurements (SF and TF) (Takahashi)
- Safety at the site (Takahashi, Ishizaki)
- No incident
2015/12/4 (Fri) 13:30 - @ Mitaka
Check the room availability (Takahashi)